### v2

# Lookahead: An Inference Acceleration Framework for Large Language Model with Lossless Generation Accuracy

Abstract.
As Large Language Models (LLMs) have made significant advancements across various tasks, such as question answering, translation, text summarization, and dialogue systems, the need for accuracy in information becomes crucial, especially for serious financial products serving billions of users like Alipay. To address this, Alipay has developed a Retrieval-Augmented Generation (RAG) system that grounds LLMs on the most accurate and up-to-date information. However, for a real-world product serving millions of users, the inference speed of LLMs becomes a critical factor compared to a mere experimental model.
Hence, this paper presents a generic framework for accelerating the inference process, resulting in a substantial increase in speed and cost reduction for our RAG system, with lossless generation accuracy. In the traditional inference process, each token is generated sequentially by the LLM, leading to a time consumption proportional to the number of generated tokens. To enhance this process, our framework, named lookahead, introduces a multi-branch strategy. Instead of generating a single token at a time, we propose a Trie-based Retrieval (TR) process that enables the generation of multiple branches simultaneously, each of which is a sequence of tokens. Subsequently, for each branch, a Verification and Accept (VA) process is performed to identify the longest correct sub-sequence as the final output. Our strategy offers two distinct advantages: (1) it guarantees absolute correctness of the output, avoiding any approximation algorithms, and (2) the worst-case performance of our approach is equivalent to the conventional process. We conduct extensive experiments to demonstrate the significant improvements achieved by applying our inference acceleration framework. Code is avaliable: https://github.com/alipay/PainlessInferenceAcceleration. footnotetext: Preprint. Under review.
Large Language Model, Lossless generation accuracy, Inference Framework, Trie tree, Single-branch draft, Multi-branch draft
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: World Wide Web Conference; May 13-17, 2024; Singaporeprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Large language models (LLMs) based on transformer architecture have exhibited remarkable performance across various benchmarks, rendering them highly valuable in numerous industries. However, despite their significant achievements in language-based tasks, LLMs still face challenges in terms of inference latency when employed in generative tasks. This drawback becomes particularly apparent in scenarios where step-wise decoding is implemented. For instance, with the GLM [1] 10B model using a single Nvidia A100 GPU, the inference speed’s theoretical limit through step-wise decoding is merely 101 tokens per second.
To further gain deeper insights into the factors affecting the LLMs’ inference latency, we conduct a comprehensive theoretical analysis focused on a specific instance. It shows that the input-output (IO) and computation consumption time for inferring each token through step-wise decoding is about 10ms and 0.07ms, respectively [^1] Memory bandwidth of A100 is 2039 GB/s, a 10B model with float16 need more than 20 GB IO, the IO time is about 10ms (20/2039), while A100 has 312 TFLOPS for float16 datatype, and forward a token need about 20B FLOPs, the calculation time is about 0.06ms (20/312000). Therefore, rather than the computation consumption time that is related to the hardware’s FLOPs capacity, the IO consumption time that is highly associated with the model’s size and the memory bandwidth is the key that significantly affects the LLMs’ overall inference latency.
Various techniques, such as quantization [2, 3], sparsity [4, 5], pruning [6, 7], distilling [8, 9], and tensor decomposition [10, 11], have been proposed to reduce the LLMs’ size and the IO consumption time for predicting each token in LLMs. However, these techniques have been found to result in a degradation of accuracy. To address the challenge of predicting more tokens within a fixed IO consumption time, non-autoregressive neural machine translation (NAT) [12] and iterative parallel decoding [13] have been introduced and successfully applied in translation tasks. Unfortunately, this approach has shown limited effectiveness in question-answering scenarios.
Recently, decoding with a draft model has become a popular strategy, as discussed in [14, 15]. However, this strategy necessitates additional training efforts or the search for an available draft model. In light of this, a strategy named LLMA that eliminates the need for a draft model has been developed, as presented in [16]. This strategy involves matching text spans from retrieved documents without requiring significant additional efforts. Despite these merits, it has been observed that this strategy performs poorly beyond document-retrieving scenarios.
Table 1 presents a summary of the aforementioned acceleration methods. It is important to acknowledge that these methods only make use of tokens in a single draft manner, thereby not fully leveraging the computational capabilities of GPUs in parallel. In order to maximize the benefits offered by GPUs, we have developed Lookahead, a groundbreaking framework that incorporates a multi-branch strategy. Lookahead utilizes a trie tree to meticulously record the input and output token lists, allowing for the prediction of multiple paths based on the provided context tokens through a parallel Verification and Accept (VA) process. Furthermore, we have implemented an adaptive strategy to optimize the path-retrieving process, effectively striking a balance between memory and computation requirements. It shows that Lookahead proves its superior performance in accelerating LLMs’ inference, compared with the existing state-of-the-art (SOTA) acceleration method.
Table 1. Comparison of different acceleration methods.
Methods Training- Accuracy- Full-FLOPs-
free lossless utility
Quantization [2, 3] - $\times$ $\times$
Sparsity [4, 5] - $\times$ $\times$
Distilling [8, 9] $\times$ $\times$ $\times$
Tensor-decomposition [10, 11] - $\times$ $\times$
Early-exit [17, 18] $\times$ $\times$ $\times$
Block decoding [14] $\times$ $\times$
Speculative decoding [15, 19] - $\times$
LLMA [16] $\times$
Lookahead
Being benefited from the superior performance and accessibility, our Lookahead framework has been widely employed in dozens of the real world scenarios of Alipay, including financial RAG, health suggestion, medical report summary, etc.
Moreover, to gain a wide range of applications, we have implemented our framework based on the transformers library of Hugging face [^2] https://huggingface.co/, by extending a generation mode named lookahead generation, which supports the greedy search and sample strategy with batch inference supported. We have also currently applied Lookahead to the most recent LLMs, such as GLM[1], Llama[20], OPT[21], GPT2[22], BLOOM[23], ChatGLM[24], Baichuan[25] and Qwen[26], etc. The aforementioned models can be easily adapted to integrate Lookahead, our well-designed framework, with only minor code modifications of approximately 20 lines. We will soon release our work with open source.
Our contributions can be summarized as:
  • We empirically quantify that the main bottleneck of LLM inference is the IO bandwidth, rather than the FLOPs.
  • We innovatively develop Lookahead, a framework that applies a hierarchical multi-branch draft strategy implemented with a trie tree to output more tokens per step than the traditional methods, in accelerating LLMs’ inference.
  • We extensively conduct experiments on both the industry and open source datasets and prove that Lookahead brings a significant improvement over the existing SOTA method in accelerating LLMs’ inference.
  • We elaborately adapt Lookahead to the most recent LLMs and will soon release our work with open source.
2. Related Work
Recently, several strategies have been proposed and developed to enhance the inference speed of LLMs while maintaining the output quality within an acceptable range. One such strategy is the non-auto-regressive approach, specifically non-auto-regressive translation (NAT) [12], primarily employed in translation tasks [27, 28, 29]. However, it is essential to note that there are significant distinctions between translation tasks and general language model (LLM) scenarios, which may lead to subpar performance when applying the NAT strategy to LLM decoding.
To address this limitation, Huang et al. [28] introduce a layer-wise iterative method wherein each layer utilizes the decoding results and embeddings from the preceding layers. This approach involves training each layer through maximum likelihood estimation to predict the outcomes of subsequent decoding layers. On the other hand, Santilli et al. [13] formalized the standard greedy auto-regressive decoding strategy by employing a parallel Jacobi and Gauss-Seidel fixed-point iteration. It initializes the next tokens using special tokens and performs iterative decoding until convergence. However, all these methods may suffer from the risk of accuracy degeneration, since the manipulated model deviates from its original version.
Recently, there has been a proposal for accuracy-lossless acceleration to enhance the auto-regressive decoding strategy as illustrated in Table 1. In this approach, a block-wise parallel decoding strategy was introduced by [14]. In this strategy, each subsequent token is independently and parallelly predicted as a proposal using an additional transformer model, which consists of a multi-output feed-forward layer and can be fine-tuned or distilled for optimal performance. Then the proposals are directly compared against the output tokens generated by the original decoder. The longest verified tokens are then selected for acceptance as the current step’s output. Despite this strategy’s remarkable acceleration, more efforts are needed to train the extra feed-forward layer, which is essential to ensure that the predictions generated by the transformer model are accurate and reliable. On the other hand, due to the parallel mechanism employed in this strategy, the prediction process solely relies on the input tokens and does not consider the information of the next token, undermining the prediction’s accuracy. Furthermore, frequent failures during the verification process shorten the output length for each decoding step.
To address the aforementioned issue, speculative decoding has been proposed [15, 19], through utilizing a smaller model as a draft predictor. For instance, the Bloom model, which has a size of 7.1B, can be employed as a small model for the larger 176B model. However, obtaining a qualified small model often becomes impractical due to the difficulty of generating similar sentence pieces as the large model. To overcome this challenge, model-free prediction strategies have been introduced to achieve accurate predictions without relying on a specific model. One such strategy is presented in Ge et al. [30], which utilizes an input-guided method that copies content from the input sentence through prefix matching. Another strategy LLMA, proposed by Yang et al. [16], employs a prefix matching approach to retrieve content from either the input sentence or a document database. These model-free strategies are utilized in a greedy decoding manner, ensuring accuracy while achieving performance gains through approximate inference. This inference technique relaxes the verification criterion by accepting predictions within the top-k scores [14]. However, it is worth noting that all the methods mentioned above utilize tokens in a single draft manner, failing to fully utilize the GPUs.
3. Preliminary
3.1. Inference Speed
The inference speed V can be expressed as below.
(1) $V=\frac{L}{T}\propto\frac{L}{N\times t(l)}$
$N\propto\frac{L}{l}$
Here, L denotes the overall generation tokens’ length and T is the overall inference time, which is positively correlative to the overall consuming time for decoding: N indicates the overall decoding steps, $t(l)$ is the decoding time per step. It should be noted that $t(l)$ is nearly constant while l, namely the generated tokens’ length per decoding step, is within a certain range, whose details and explanations will be introduced and discussed in the following sub-section. Therefore, given the fixed L, the longer l is, the fewer N is needed, which in turn promises a higher inference speed V.
3.2. Step-Wise Decoding
Auto-regressive language models have been firstly introduced through following a step-wise decoding strategy: at each decoding step, the models concatenate the prompt sequence and the previous generated tokens and output the next single token using greedy-decoding, which selects the next token with the highest predicted probability over the vocabulary.
Though this strategy has been widely applied, the particular process promises only one single output token per decoding step ($l=1$), which limits the overall inference speed.
3.3. Singe-Branch Strategy
Several most recent methodologies [14] [16] have been proposed to generate a sequence of tokens at each decoding step, with the purpose of promising a higher l to accelerate the LLMs’ inference speed V. In these works, according to the prompt and the output tokens at the previous decoding step, a branch of tokens, named single-branch draft, have been obtained through the small model or the document reference, and efficiently validated by running the LLM in a single forward process.
Despite the single-branch strategy’s success in accelerating LLMs’ inference, the generated tokens’ length per decoding step, namely l, cannot be guaranteed as long as we wish. The previous works [14] [16] empirically conclude that there is an upper limit for l and the inference speed V.
This can be explained through the successive validating mechanism: the validation process breaks once one token in the single-branch draft fails the validation, only the validated tokens in front of this failed token are accepted as the output. For the sake of brevity, we give a definition below:
Definition 1 Following the predicted next token, only the successive validated tokens from the beginning of the branch draft are kept and accepted as the output tokens, whose length is called the effective decoding length, EDL.
Simply extending the single-branch draft’s length over EDL not only fails to promise a longer output tokens’ length per decoding step l and a higher inference speed V, but also wastes more computation. Apparently, how to achieve a longer EDL, namely the effective decoding length, is the key to further accelerate LLMs’ inference.
3.4. Discovering GPU’s FLOPs Redundancy
To explore how far EDL can be extended at each decoding step, we conduct a case study through discussing the decoding length’s impact on the overall consuming time of LLMs’ single forward process, whose result can be found in Figure 1. It should be noted that we apply AntGLM-10B (which is trained from scratch with the GLM structure [1, 24]) model with single Nvidia A100 GPU as an example.
Figure 1. Decoding length’s impact on the overall consuming time of LLMs’ single forward process.
In Figure 1, it is evident that as the decoding length increases in the LLM’s single forward process, the time consumption remains relatively constant in the initial stage, given a fixed context length. This observation can be attributed to several factors. Firstly, certain time requirements, such as kernel launching, operator loading, reading, and verification, are fixed and dependent on the GPU bandwidth. These factors are unrelated to the computational complexity or decoding length. Secondly, when the decoding length is relatively short, the resulting small matrix blocks may not take full advantage of CUDA and Tensor cores, indicating that GPU shows its FLOPs redundancy in this scenario. Consequently, the differences in overall time consumption between decoding lengths of 16 and 128 in the LLM’s single forward process is negligible, provided the context length is 64 or 256. This finding supports the concept that the decoding time per decoding step, represented as t(l), remains nearly constant within a specific range of generated token lengths l. However, when the decoding length exceeds 128 while the context length is 64 or 256 in the LLM’s single forward process, the larger matrix block size introduces a more complex calculation logic for the tensor cores. Consequently, the overall inference time shifts to the second stage, leading to a gradual increase in overall inference time.
Definition 2 We define the maximum decoding length within which the inference consuming time of the single forward process is nearly constant as the critical decoding length, CDL.
It can be clearly seen that there is still a ”sufficient gap” between EDLs empirically concluded previously [14] [16] and the CDL illustrated in Figure 1. In this scenario, we come up with two questions:
1. How to extend EDL further to accelerate the LLMs’ inference as much as possible?
2. How to fully leverage the capabilities of GPUs, particularly considering our preliminary finding that GPUs exhibit FLOPs redundancy when the decoding length is within the CDL?
Motivated by these questions, we propose a framework called Lookahead that aims to achieve a longer EDL by optimizing GPU utilization. The subsequent section provides detailed insights into this framework.
4. Methods
4.1. Overview of Lookahead
We construct a framework, Lookahead, to accelerate LLMs’ inference through developing and applying the multi-branch strategy.
In contrast to the single-branch strategy such as [16], which only considers one available draft without considering the existence of other drafts, the multi-branch strategy retrieves multiple drafts simultaneously. These drafts are then efficiently decoded and validated in parallel through the Verification and Accept (VA) process, progressively. The VA process then identifies the correct sub-sequence for each draft and retains the longest sub-sequence as the output tokens. Figure 2 presents the progress overview including various draft retrieving and verifying strategies.
The execution of the multi-branch strategy is based on a fundamental fact that when considering a sequence of tokens, there may exist multiple sequences of successive following tokens, each representing a potential branch draft. These sequences are referred to as multi-branch draft in our work, whose details will be introduced in the next sub-section.
Figure 2. Overview of the drafts retrieving and the Verification and Accept (VA) process using various strategies.
4.2. Multi-Branch Draft
While developing and applying the multi-branch strategy, a concern naturally arises regarding how to efficiently deal with as many drafts as possible within the constraints of limited computational resources, specifically the CDL.
Fortunately, through careful observation of the context, we have noticed that certain branches in the drafts have common prefix tokens. For example, in Figure 2, the branches [on, my, knee] and [on, a, table] both share the prefix token [on]. By leveraging this observation, we can compress and organize the multi-branch draft using a hierarchical structure, allowing for the inclusion of additional branches while maintaining the same token capacity. To achieve this, we recursively merge the prefix token(s) shared by multiple branches, and then append the remaining tokens of each branch. We refer to these compressed and organized drafts as hierarchical multi-branch draft. On the other hand, drafts organized in a parallel manner without using the hierarchical structure are referred to as parallel multi-branch draft.
As illustrated in Figure 2 describing the process of an inference step, through the single-branch strategy, the single draft [on, my, knee] is appended to the next token [sits] directly for the single forward process in VA, which in turn gives the output [on, a]. Being benefited from the multi-branch strategy with the parallel multi-branch draft, two branches [on, my, knee] and [on, a, table] are organized in parallel and appended to the next token [sits] for verifying, which outputs [on, a, chair]. By utilizing the hierarchical multi-branch draft, the shared prefix token [on] are merged, after which the token list [my, knee, a, table] are appended. By doing so, given the same decoding length, we are able to save an additional space to append one more token [chair] from another branch [on, a, chair]. Combined with the next token, we collect the token list [sits, on, my, knee, a, table, chair] for decoding through the single forward process and output [on, a, chair, and]. Therefore, compared to the single-branch strategy, our multi-branch strategy, particularly when employing the hierarchical multi-branch draft, offers the advantage of retrieving multiple drafts, resulting in improved EDL from a statistical perspective and thus significantly enhancing LLMs’ inference speed.
The position IDs and causal masks in a transformer-based model [31] are also merged to align with the merged token list. By doing so, we are able to accommodate more branches compared to the utilization of parallel multi-branch draft, whose details are presented in Figure 3.
To facilitate the hierarchical structure of branches, we utilize a trie tree, a data structure commonly used for efficient retrieval and storage purposes to record more candidate branches, whose details will be introduced in the next section.
Figure 3. The input ids, position ids and causal masks for forwarding using various strategies.
4.3. Trie Tree
The trie tree is a data structure that efficiently handles prefix matching by organizing nodes as individual characters or words. In our work, each node of the trie tree represents a token ID, and a path from the root to a leaf node represents a branch.
Before and after each step consisting of the draft retrieving process and the VA process as illustrated in Figure 2, a global trie tree will be updated through multiple procedures, which will be introduced in the sub-section 4.3.1. During the draft retrieving process, the trie tree will be retrieved to provide the drafts, whose details will be introduced in the sub-section 4.3.2.
4.3.1. Trie Tree Updating
We apply the branch inserting, eliminating and node pruning to update the trie tree, in order to maintain the trie tree’s size and keep the most frequency tokens.
Branch Inserting. Considering that in several typical scenarios (e.g., RAG, Summary), the output are often derived from the prompt, we thus manipulate the prompt into the branches and insert them into the trie tree. Moreover, towards various prompts, there may be similar outputs, e.g., ”keep exercising” in answering ”how to keep healthy” and ”how to be stronger”. Therefore, we also insert the outputs into the trie tree. Given the prompt or the output, each sub-sequence of the successive tokens with the specified length can be regarded as a branch draft and this specified length is the branch length. Each branch is then automatically inserted into the trie tree, whose construction process has been widely applied in document matching [32, 33].
Branch Eliminating. When the current generation in answering the prompt is finished after multiple steps, the branches derived from this prompt are eliminated, considering that these branches may not be relevant to the generation in answering the other prompts.
Node Pruning. To maintain the trie tree within a moderate size, in case that the oversized trie tree results in high memory consumption and slow retrieval performance, we develop a pruning strategy. To achieve this, we remove the least frequent nodes dynamically, when the trie tree exceeds a predetermined capacity. By doing so, we are able to optimize memory consumption and improve retrieval performance in our framework.
The analysis of the above updating procedures will be discussed in the experiment section.
4.3.2. Trie Tree Retrieving
We can extract multiple branches from a constructed trie tree by providing a prefix, which is essentially a list of tokens. The length of the prefix influences the number and relevance of the retrieved branches. Short prefixes yield a greater number of branches, while longer prefixes result in more related branches. To strike a balance between the count and correlation of the branches, we adopt a multi-stage retrieval strategy inspired by [16]. Specifically, we begin by attempting to match a longer prefix. If the number of tokens associated with the matched branches is significantly smaller than the CDL, we reduce the length of the prefix and retry the matching process until we obtain a substantial number of tokens linked to matched branches.
If the count of the matched branches falls below a predefined threshold, we utilize all of them for the VA process. However, if the count exceeds a given size, we choose the tokens with the highest frequency. To prioritize the importance of branches derived from the input prompt, we amplify their frequency by a significant factor.
4.4. Work Flow
We summarize the work flow of Lookahead with pseudo code in the Algorithm 1.
Algorithm 1 Lookahead with multi-branch draft
0:  decoding length $L_{d}$, branch length $L_{b}$, minimal count without re-retrieving $N_{m}$, node capacity of trie tree $N_{max}$, start token id $S$
0:  $O$
1:  initialization: trie tree $\mathcal{T}$
2:  repeat
3:     initialization: output id list $O$=[$S$], KV cache $C_{kv}=[]$
4:     tokenize a query to token list $T$
5:     for $i\leftarrow 1$ to len($T$)-1 do $\triangleright$ branch inserting
6:        insert $T[i:i+L_{b}]$ into $\mathcal{T}$
7:        if Node count of $T$ $N_{m}$  then
8:           do node pruning
9:        end if
10:     end for
11:     repeat
12:        for $j\leftarrow len(O)$ to 1 do $\triangleright$ trie tree retrieving
13:           obtain a prefix token list $T_{p}\leftarrow O[-j:]$
14:           match a sub trie tree $\mathcal{T}_{s}$ from $\mathcal{T}$ with $T_{p}$
15:           if Node count of $T_{s}$ $N_{m}$  then
16:              continue
17:           end if
18:           select $L_{d}$-largest frequency nodes $N_{max}$ from $\mathcal{T}_{s}$
19:        end for
20:        prepare token ids $T_{fp}$, position ids $P_{fp}$ and causal masks $M_{fp}$ for $N_{max}$
21:        get next tokens prediction via $LLM(C_{kv},T_{fp},P_{fp},M_{fp)}$
22:        verify each branches and accept verified tokens $T_{v}$
23:        append $T_{v}$ to $O$
24:        incrementally put branches of $O$ to $\mathcal{T}$
25:        rearrange KV cache $C_{kv}$ with accepted tokens
26:     until meet stopping criteria
27:     eliminate frequency of branches from the current prompt
28:  until all queries are procossed
5. Experiments
5.1. Experimental Setup
We conduct a solid experiment to evaluate the efficacy, robustness and practicality of our Lookahead framework in accelerating LLMs’ inference. Our inference environment includes a server with 32-core CPU, 64 GiB host memory, and a A100-SXM(80G) GPU. Considering the actual industry scenarios that Lookahead is applied, AntRAG is chosen as the evaluation dataset. AntRAG dataset is an internal benchmark dataset that is meticulously collected from a real-life product system in Alipay. The dataset collection process involves submitting user queries to a search engine, which then retrieves the most relevant document alongside the original query. These query-document pairs are combined to form prompts. To ensure the dataset’s quality and coherence, we employ a rigorous process of removing any duplicate or unrelated prompts, resulting in a refined and comprehensive dataset. Furthermore, to validate the robustness and practicality of Lookahead in open-domain scenarios, we additionally conduct experiments utilizing the Dolly dataset  [^3] Dolly dataset: https://huggingface.co/datasets/databricks/databricks-dolly-15k. For the evaluation purposes, considering that all the candidate methods promise the accuracy lossless generation, therefore the inference speed with the unit of output tokens per second is regarded as the unique evaluation metric.
In evaluating the applicability of base LLMs across different datasets, it is essential to consider their adaptability. For our study, we select Antglm-10B as the base model for the AntRAG dataset, while Llama2-13B serves as the base model for the Dolly dataset. Detailed information regarding the datasets and base models utilized can be found in Table 2 and Table 3, respectively. It should be noted that for each dataset, we use the test set containing 1000 samples for the inference speed evaluation, while applying the dev set for warm-up to simulate the real-life scenarios.
Table 2. Summary of datasets.
Dataset Split #samples #tokens of prompt #tokens of output
AntRAG dev 7,605 243.1 81.9
test 1,000 241.0 82.0
Dolly dev 13,850 298.3 101.8
test 1,000 301.5 104.8
Table 3. Summary of models.
Model Antglm-10B Llama2-13B
Params 10.14B 12.71B
Vocab size 115328 32000
Layer 48 40
Hidden size 4096 5120
Attention head 32 40
MLP intermediate size 16384 13824
5.2. Results
5.2.1. Lookahead’s superior performance
Table 4 exhibits the inference speeds achieved by different acceleration methods. To minimize the influence of random factors, we conduct the evaluation process five times, ensuring the accuracy and reliability. The mean value of the inference speed is then selected as the performance indicator for each acceleration method.
Table 4. Inference Speed of different methods and datasets. The results are calculated by conducting five repeated runs.
Methods AntRAG Dolly
token/s speedup token/s speedup
Baseline 52.4±0.2 1.00 34.0±0.3 1.00
LLMA 165.4±3.3 3.16 50.8±0.2 1.49
Lookahead (Parallel) 263.4±3.3 5.03 68.9±0.4 2.03
Lookahead (Hierarchical) 280.9±4.4 5.36 71.7±0.9 2.11
As depicted in Table 4, the average inference speed of AntGLM-10B towards AntRAG is recorded at 52.4 tokens/s. However, with the incorporation of LLMA, this speed is elevated to 165.4 tokens/s, resulting in a notable 3.16 times speed improvement. Implementing Lookahead with parallel multi-branch draft further enhances AntGLM-10B’s average inference speed to 263.4 tokens/s, achieving a remarkable 5.03 times speed-up, representing a significant acceleration. Additionally, the adoption of hierarchical multi-branch draft via trie tree propels the average inference speed even further, culminating in a 5.36 times speed-up, surpassing LLMA by 70% in terms of acceleration.
In the utilization of Llama2-13B as the base model for inferring the Dolly dataset, the observed speed enhancements resulting from different accelerations may not exhibit significant disparities. It should be noted that the Dolly is an open-domain dataset. In this dataset, the output tokens may not occur in the input prompts, which differs from the real-world scenarios where the output tokens often share a common topic and can be predicted more easily. Nevertheless, Lookahead consistently demonstrates its remarkable superiority and applicability across diverse datasets, further emphasizing its practicality in real-world scenarios.
5.2.2. Hyper-parameters in multi-branch draft
In continuation of AntGLM-10B, we delve deeper into the analysis of the decoding and branch lengths, two key hyper-parameters within our Lookahead framework. In particular, we empirically examine their impact on the inference speed, as illustrated in Figure 4.
Figure 4. The decoding and branch length’s impact on the LLM’s inference speed using various accelerations.
Generally speaking, it can be observed that as the decoding and branch lengths increase, there is an upward trend in the overall inference speed. As mentioned, with the single-branch strategy, LLMA is unable to further enhance the inference speed while its branch length surpasses 25. This limitation arises due to its inability to ensure a longer EDL. By implementing the multi-branch strategy, Lookahead is able to retrieve multiple branch drafts and thus guarantee a longer EDL after the VA process, as illustrated in Figure 5. Consequently, this leads to a significantly higher inference speed, while maintaining the same branch length as LLMA in Figure 4. Furthermore, the improved EDL as shown in Figure 5 also provides more headroom for increasing the branch length and thereby enhancing the upper limit of the inference speed.
In addition, Figure 5 shows that being benefited from dealing with more branch drafts, Lookahead using the hierarchical multi-branch draft promises the advantage of a longer EDL and consequently improved inference speed, compared to using the parallel multi-branch draft with the identical branch and decoding lengths. It should be noted that in Figure 5, improving the decoding length promises a longer EDL using Lookahead, however the oversized decoding length that surpasses the CDL fails to promise a higher inference speed, due to a more complex calculation logic for the tensor cores that has been introduced in the preliminary.
Figure 5. The decoding and branch length’s impact on the effective decoding length, EDL using various accelerations.
5.2.3. Hyper-parameters in trie tree updating
Table 5 showcases the inference speeds of Lookahead, which exhibit variability depending on different procedures used for updating the trie tree. These procedures include branch inserting with prompt and/or output, branch eliminating, and node pruning. In comparison to the results obtained without a specific procedure, Lookahead attains its optimal performance when all the aforementioned procedures are applied to dynamically update the trie tree, highlighting the necessity of these procedures in updating the trie tree.
Table 5. The inference speed of different updating procedures was evaluated, considering various conditions. In the evaluation, ”W/o prompt” refers to the absence of branching from the prompt to the trie tree, while ”W/o output” indicates the absence of branching from the output to the trie tree, ”W/o Pruning” signifies the exclusion of node count reduction in an oversized trie tree through pruning, ”W/o eliminating” denotes the retention of branches derived from the prompt before processing the subsequent query.
Condition W/o prompt W/o output W/o pruning W/o elimin Lookahead
Token/s 234.5 202.0 221.4 234.2 280.9
The capacity of the trie tree has a significant impact on the acceleration performance. To determine the optimal capacity, we conduct empirical experiments, as outlined in Table 6. Instead of using a fixed value, we normalize the node capacity in relation to the decoding length, which allows us to identify a suitable hyper-parameter that performs well across different decoding lengths. Based on our findings, we conclude that node capacities of 16 and 32 times the decoding length yield the best results. Consequently, we have set the default capacity in our framework to be 16 times the decoding length.
Table 6. Inference Speed of different capacity of trie tree. n*DL denotes the capicity is n times the decoding length.
Capacity 1*DL 2*DL 4*DL 8*DL 16*DL 32*DL 64*DL 128*DL
Token/s 254.5 258.7 268.8 273.3 280.7 280.8 279.5 278.3
5.2.4. Lookahead’s efficient memory usage
Figure 6 shows the peak GPU memory consumption for various decoding lengths (1, 2, 4, 8, 16, 32, 64, 128). The decoding length of 1 represents the experience where no Lookahead is used, resulting in a GPU memory consumption of 20.248 GiB. Surprisingly, even with a decoding length of 128, the use of Lookahead only leads to a negligible increase in GPU memory consumption of 0.6%. This minimal increase in memory usage is practically negligible and has no significant impact on real-world applications. Furthermore, we have also evaluated the CPU memory consumption of the trie tree. With the AntRAG dataset, the trie tree only utilizes a mere 260 MiB of memory. This value is negligible when compared to the total memory capacity of a mainstream server.
Figure 6. GPU Memory (i.e., device memory) of different decoding length. The GPU memory denotes the peak memory consumption in a inference process.
5.3. Online Deployment
Our framework has been widely used in real-world scenarios of Alipay since April 2023, due to its accessibility, accuracy-lossless and remarkable acceleration. To be concise, we only report 5 typical scenarios in Table 7. CHART2JSON is a scenario that uses a multi-modal model to convert a chart image to structured content with JSON format, it can achieve extraordinary acceleration due to plenty of template tokens in generated JSONs. Citizen biz agent, enterprise info QA and health suggestion are RAG scenarios, they are used for answering user questions with reference documents. Medical report summary is a scenario to summarize the content (texts after OCR) of a medical report image. With the assistance of our lookahead framework, the latency of all scenarios is decreased significantly while the generation results are the same as the original step-by-step strategy.
Table 7. Inference Speed of product scenarios. The speed is measured by mean latency (seconds) of a query.
Scenario Baseline Lookahead Speedup
CHART2JSON 12.97 2.07 6.26x
Citizen Biz Agent 1.67 0.32 5.21x
Enterprise Info QA 14.01 2.74 5.11x
Health Suggestion 12.41 2.66 4.66x
Medical Report Summary 3.33 1.25 2.66x
6. Conclusion
In our work, we empirically quantify that the main bottleneck of LLM inference is the IO bandwidth, rather than the FLOPs. Inspired by this, to take full advantage of the GPU’s FLOPs redundancy, we innovatively develop Lookahead, a generic framework that applies a hierarchical multi-branch draft strategy implemented with a trie tree to output more tokens per step than the traditional methods. We conduct extensive experiments and demonstrate that Lookahead gains a substantial inference acceleration and cost reduction, with lossless generation accuracy. By simply adapting our framework to the latest LLMs, we have achieved a wide range of applications and promising prospects and will soon release our work with open source.
Appendix A Batch Inference
It is important to note that the aforementioned experiments were conducted with a batch size of 1. However, as illustrated in Table 8, Lookahead also demonstrates its effectiveness with non-singular batch sizes, achieving significantly higher inference speeds compared to step-by-step batch inference and LLMA-based batch inference strategies. Although the speedup is not as remarkable as with singular batch sizes, there are two key reasons for this. Firstly, as the batch size increases, the need for parallel inference begins to strain the available redundancy of the GPU. This can limit the potential speedup that can be achieved. Secondly, different samples within a batch are at different stages of progress, meaning that some may have generated fewer tokens compared to others. This variability in progress leads to additional time consumption in attention calculation, further impacting the overall inference speed. However, it is important to highlight that Lookahead is primarily focused on time-sensitive scenarios rather than throughput-sensitive scenarios. In the future, we plan to optimize the throughput by implementing continuous batching and paged attention mechanisms [34].
Table 8. Inference Speed of different batch sizes with the AntRAG dataset.
Method Batch size=2 Batch size=4
token/s speedup token/s speedup
Baseline 68.0 1.00x 88.2 1.00x
LLMA 185.2 2.72x 214.0 2.43x
Lookahead(Parallel) 274.4 4.04x 291.2 3.30x
Lookahead(Hierarchical) 285.5 4.20x 299.5 3.40x
Appendix B Prompt templates with the Dolly dataset and Llama-13B model.
Similar to [35], we construct a prompt with the below template.
Table 9. Prompt templates of Dolly.
Prompt Type Propmt Template
W/ reference
Below is an instruction that describes a task, paired
with an input that provides further context. Write
a response that appropriately completes the request.
### Instruction: {instruction}
### Input: {reference}
### Response:
W/o reference
Below is an instruction that describes a task. Write
a response that appropriately completes the request.
### Instruction: {instruction}
### Response: